The increase in the volume of 10cm3 of mercury when the temperature rises by 100oc is 0.182cm3. What is the cubic expansivity of mercury?


1
Expert's answer
2022-01-31T07:34:24-0500

Given:

"V_0=10\:\rm cm^3"

"\Delta V=0.182\:\rm cm^3"

"\Delta T=100\:\rm^\circ C"


The volume thermal expansion in described by an equation

"\Delta V=V_0\gamma\Delta T"

Hence, the cubic expansivity

"\gamma=\frac{\Delta V}{V_0\Delta T}=\frac{0.182}{10*100}=1.82*10^{-4}\:\rm K^{-1}"